OpenStudy (anonymous):

Given tan(θ) = - 2⁄3 and π⁄2 < θ < π , find sin(θ).

OpenStudy (anonymous):

Well tan=O/A. So you have a right triangle. The hypotenuse is (2)^2+(-3)^2=c^2=sqrt(13). So cos is negative in the second quadrant (which is pi/2 to pi) So its positive: sin(theta)=2/sqrt(13)

OpenStudy (anonymous):

thanks, the denom needed radical factorized again - that;s where i missed it

OpenStudy (anonymous):

(2/13)(sqrt13)

OpenStudy (anonymous):

Haha, nice :)